# Web Archives Metadata Generation with gpt-4o: Challenges and Insights

### v1

@@ 4.1 Data Reduction
From the DataFrame, each heuristic finished within 15 seconds on 112 files. The heuristics were consolidated. Based on our heuristics, the final DataFrame should have an identical number of rows as the number of WARC files. Based on the collection of 112 WARC files we have, we obtain a 99.9% reduction in total token count, and consequently costs compared to letting OpenAI parse the entirety of the WARC files.
Figure 4: Visualizing cost differences before and after applying heuristics to count tokens with tiktoken [39], a Byte-Pair Encoding tokenizer for gpt-4o.
@@ 4.2 Statistical Test
We performed Cochran’s Q test by having trained cataloguers rate the titles and abstracts, without knowledge of their provenance. Based on a 5% confidence level, we found that the synthetic titles and abstracts are statistically distinguishable from human-generated titles and abstracts. Since the p-value (0.02) is less than our significance level ($\alpha=0.05$), we reject the null hypothesis, indicating a significant difference between the heuristic-generated and human-generated metadata.
@@ 5.1 Implications of Results
    Quality Comparison: Our adaptation of the Turing Test, involving 8 cataloguers evaluating metadata from different sources, provided valuable insights into the current capabilities of LLMs, specifically gpt-4o in this domain. The synthetic titles and abstracts were statistically distinguishable from human curated metadata, with a p-value of 0.02, suggesting that human generated metadata still maintains an edge in quality. Furthermore, there is no significant difference between LLM based approaches, with rules and without rules.
@@ 5.2 Limitations and Challenges
    Quality of Input Data: The adage "garbage in, garbage out" [42, 43] is especially relevant in our context. The quality of LLM generated metadata is directly dependent on the quality of the input data from web archives. Ensuring that our web archive data is accurate, diverse, well labelled, and free from noise is crucial for generating reliable and unbiased metadata.
@@ 6 Conclusion
Our study marks a significant step towards leveraging AI in web archiving, offering valuable insights into its current capabilities and limitations. Addressing the identified challenges will help us work towards a future where AI enhances the preservation and accessibility of digital heritage, maintaining high standards crucial for the utility of web archives. This research sets a roadmap for future improvements, aiming to bridge the quality gap between AI and human curated metadata.

### v3

@@ 4.1 Data Reduction
From the DataFrame, each heuristic finished within 15 seconds on 112 files. The heuristics were consolidated. Based on our heuristics, the final DataFrame should have an identical number of rows as the number of WARC files. Based on the collection of 112 WARC files we have, we obtained a 99.9% reduction in total token count, resulting in significant cost savings compared to letting OpenAI parse the entirety of the WARC files.
Figure 4: Visualizing cost differences before and after applying heuristics to count tokens with tiktoken, a Byte-Pair Encoding tokenizer for GPT-4o [39].
@@ 4.2 Statistical Test
We performed Cochran’s Q test by having trained cataloguers rate the titles and abstracts, without knowledge of their provenance. Based on a 5% confidence level, we found that the synthetic titles and abstracts are statistically distinguishable from human-generated titles and abstracts. Since the p-value (0.02) is less than our significance level ($\alpha=0.05$), we reject the null hypothesis, indicating a significant difference between the GPT-4o generated metadata and human crafted metadata.
@@ 5.1 Implications of Results
    Quality comparison: Our adaptation of the Turing test, involving eight cataloguers evaluating metadata from different sources, provided valuable insights into the current capabilities of LLMs, specifically GPT-4o in this domain. The synthetic titles and abstracts were statistically distinguishable from human curated metadata, with a p-value of 0.02, suggesting that human generated metadata still maintains an edge in quality. Furthermore, there is no significant difference between LLM based approaches, with rules and without rules.
@@ 5.2 Limitations and Challenges
    Quality of input data: The adage "garbage in, garbage out" is especially relevant in our context. The quality of LLM generated metadata is directly dependent on the quality of the input data from web archives. Ensuring that our web archive data is accurate, diverse, well labelled, and free from noise is crucial for generating reliable and unbiased metadata [42, 43].
@@ 6 Conclusion
Our study marks a significant step towards leveraging AI in web archiving, offering valuable insights into its current capabilities and limitations. Addressing the identified challenges will help us work towards a future where AI enhances the preservation and accessibility of digital heritage, maintaining high standards crucial for the utility of web archives. This research sets a roadmap for future improvements, aiming to bridge the quality gap between AI and human curated metadata. Our study marks a significant step towards leveraging AI in web archiving, offering valuable insights into its current capabilities and limitations. Addressing the identified challenges will help us work towards a future where AI enhances the preservation and accessibility of digital heritage, maintaining high standards crucial for the utility of web archives. This research sets a roadmap for future improvements, aiming to bridge the quality gap between AI and human curated metadata.